OpenStudy (anonymous):

a curve on a set of railroad tracks consists of 2 circular arcs. The central angle of one arc is 20 with a radius of 2500m. The central angle of the 2nd arc is 25 with a radius of 3000m. find the combined length of the 2 arcs to the nearest meter

OpenStudy (anonymous):

Formula to find an arc's length : S = R a R: radius ; a central angle. Combined length: S = R1*a1+ R2*a2 S = 20*2,500 + 25*3,000 = ...